Make a study about how many sheets of paper you consumed weekly in

answering your Self Learning Modules. Record the quantity (total number of sheets)

per subject, then construct a probability distribution. Compute the mean, variance,

and the standard deviation of the probability distribution you made. Interpret the

result, then find out how many weeks you will consume 50 sheets of pad paper.


1
Expert's answer
2021-04-26T04:06:31-0400


Mean number of sheets "\mu = \sum X. P(X)=0.15+0.56+0.42+1.04+0.85=3.02"


Variance "\sigma^2=" "\sum[X^2.P(X)]-\mu^2= [0.15+1.12+1.26+4.16+4.25]-3.02"

"=10.94-3.02=7.92"


Standard Deviation "\sigma = \sqrt{7.92}=2.81"


So, no. of weeks required to use 50 sheets of paper = "\dfrac{50}{3.02}=16.56 \approx 17 \ weeks"